Hira has Rs.1200 a week to spend on yogurt and mangoes. The price of yogurt is Rs.200, and mangoes are Rs.400 a box. (Hint: Quantity of mangoes to be plotted on the x-axis of the budget line)

  1. Calculate the Y-intercept for the budget line. 
  2. Calculate the X-intercept for the budget line. 
  3. Calculate the slope of the budget line. 
1
Expert's answer
2021-06-02T10:43:46-0400

Solution:

1.). The Y-intercept for the budget line ="\frac{M}{Py}"

M = Total income = Rs.1,200

Py = Price of yogurt = Rs.200


The Y-intercept = "\frac{1,200}{200} = 6"


The Y-intercept = 6

 

2.). The X-intercept for the budget line = "\frac{M}{Px}"

M = Total income = Rs.1,200

Px = Price of mangoes = Rs.400


The X-intercept = "\frac{1,200}{400} = 3"


The X-intercept = 3

 

3.). The slope of the budget line is the ratio of the prices of good x and y (mangoes and yogurt). This means the price of the good on the x-axis divided by the price of the good on the y axis. The slope of the budget line is always negative as it is downward sloping.


The slope of the budget line = "\frac{-Px}{Py} = \frac{-400}{200} = -2"


The slope of the budget line = -2